### v1

# The Manakov-Zakharov-Ward model as an integrable decoupling limit of the membrane

Abstract
[…]
1 Introduction
[…]
In addition to these previously mentioned understood corners of membrane theory, this letter suggests a novel tractable regime – via a decoupling limit of the membrane in a certain background geometry – described by a $(1+2)$-dimensional classically integrable theory. Integrable field theories in $(1+1)$ dimensions evade typical no-go constraints by admitting infinitely many local conserved charges that render scattering purely elastic and factorised. However, in $(1+2)$ or higher dimensions, the same infinite symmetries force the S-matrix to be trivial (that is, free), as any non-trivial scattering would violate the independence from impact parameters. Thus, beyond one spatial dimension, integrable models must break Lorentz invariance. One of the few existing examples of such non-Lorentz invariant integrable theories is the Manakov-Zakharov-Ward model [7, 8].
A non-relativistic decoupling limit, as proposed here, fits into a recently observed pattern that certain decoupling limits that isolate BPS $p$‑brane dynamics and associated matrix models [9, 10, 11], naturally giving rise to non‑Lorentzian target‑space geometries known as $p$-brane Newton-Cartan geometry (see [12] for a recent review). More recently, this has also been discussed extended to eleven-dimensional supergravity [13, 14, 15, 16].
1.1 The bosonic membrane
The action of the bosonic membrane (or equivalently the bosonic part of the supermembrane [2]) is
$S$ $\sim\int\mathrm{d}^{3}\sigma\sqrt{\gamma}\left(\frac{1}{2}\gamma^{MN}\partial_{M}X^{\mu}\partial_{N}X^{\nu}g_{\mu\nu}\right.$ (1.1)
${}\qquad+\left.\frac{1}{3!}\epsilon^{KLM}\partial_{K}X^{\mu}\partial_{L}X^{\nu}\partial_{M}X^{\rho}C_{\mu\nu\rho}+\frac{1}{2}\right)$
in a bosonic supergravity background described by a metric $g$ and three-form gauge field $C$. $\sigma^{M}$ are the coordinates on the membrane world-volume, $X^{\mu}(\sigma)$ the embedding coordinate fields, and $\gamma_{MN}$ is the (dynamical) world-volume metric. The cosmological constant term, as usual, ensures that the world-volume metric becomes the induced metric (as algebraic equation of motion for the world-volume metric $\gamma$):
$\gamma_{MN}=\partial_{M}X^{\mu}\partial_{N}X^{\nu}g_{\mu\nu}.$ (1.2)
In contrast to the string, reparameterisation invariance is not sufficient to gauge-fix it to be the flat metric. The equations of motion for the world-volume coordinate fields $X^{\mu}(\sigma)$ are
00 0 $=\partial_{M}(\gamma^{MN}\partial_{N}X^{\mu})+\gamma^{MN}\Gamma^{\mu}{}_{\kappa\lambda}\partial_{M}X^{\kappa}\partial_{N}X^{\lambda},$ (1.3)
${}\quad+\epsilon^{NKL}\frac{1}{3!}{F^{\mu}}_{\nu\kappa\lambda}\partial_{N}X^{\nu}\partial_{K}X^{\kappa}\partial_{L}X^{\lambda},$
where $\Gamma$ denote Christoffel symbols and $F=\mathrm{d}C$. Due to the non-trivial nature of the world-volume metric $\gamma$, these equations are quite intractable, in general.
Famously [17, 18], even this supposedly most simple case leads to complications, such as a continuous spectrum and spike instabilities. Quantum non-integrability of the membrane in the sense of scattering of fluctuations around a classical solution, and its relation to string theory, has recently been shown in [19]. In this letter, we will see that a large wrapping limit together with a scaling limit of the background (or alternatively: a non-relativistic limit of fluctuations around a wrapped membrane) can simplify these equations. In the special case considered here, the result will even be a classically integrable system.
1.2 The Manakov-Zakharov-Ward model
Classically integrable field theories in dimensions higher than $(1+1)$ are rare, especially those of the $\sigma$-model type, which are abundant in $(1+1)$ dimensions (see [20, 21] for recent reviews). A notable exception is the Manakov-Zakharov-Ward (MZW) model [7, 8], an integrable field theory in $(1+2)$ dimensions. It arises via dimensional reduction from the anti-self-dual Yang-Mills equations in $(2+2)$-dimensional spacetime. This reduction involves assuming that the gauge potentials are independent of one spatial coordinate, resulting in a non-Lorentz invariant system in $(1+2)$ dimensions.
The model is of $\sigma$-model type and can be formulated in terms of maps $g:\Sigma\rightarrow G$ from a $(1+2)$-dimensional flat spacetime $\Sigma$ (referred to as the world-volume) into a Lie group $G$. The dynamics is conveniently expressed using the left-invariant Maurer–Cartan current $j=g^{-1}\mathrm{d}g\in\mathfrak{g}$, valued in the Lie algebra $\mathfrak{g}$ of $G$. The flatness condition $\mathrm{d}j+\frac{1}{2}[j,j]=0$ is automatically satisfied. The equations of motion take the form
$\mathrm{d}\star j+\mathrm{d}j\wedge v=\mathrm{d}\star j-\frac{1}{2}[j,j]\wedge v=0,$ (1.4)
where $v$ is a constant world-volume vector that explicitly breaks Lorentz invariance in $(1+2)$ dimensions.
Originally introduced by Manakov and Zakharov [7] for the case of $G=\mathrm{SU}(2)$, the model was shown to be solvable by the inverse scattering transform. It supports soliton solutions localised in all spatial directions and does not show classical scattering, an identifying feature of integrability [8]. Ward [8] further demonstrated that for the model to have a well-defined energy functional, the vector $v$ must be space-like, although its integrability remains unaffected by the nature of $v$. The MZW-model possesses an infinite hierarchy of conserved charges arising from its Lax pair and zero-curvature formulation [8]. Although only the energy-momentum tensor yields local, finite Noether charges, Ioannidou and Ward constructed an infinite sequence of non-local but finite conserved charges using integrals of the currents and its powers [22]. These charges reflect the underlying loop-algebra and twistor-theoretic structure of the model [23], and parallel the hierarchy of integrable flows found in systems such as the matrix KP-hierarchy [24]. This integrable structure also governs soliton interactions and scattering phenomena [25].
Recently, the MZW-model has reemerged in the context of higher-dimensional Chern–Simons theory [26, 27], a unifying framework for integrable $\sigma$-models associated with Lie group symmetries. In this setting, the MZW-model appears as a special case in certain higher-dimensional generalisations [28, 29]. Partly because of its breaking of Lorentz invariance, it was not obvious in which physical settings the MZW-model appears. Moreover, it seemed unlikely that the MZW-model will play a role in membrane dynamics since it assumes a flat metric on $\Sigma$ in contrast to the dynamical induced metric of the bosonic membrane. However, this article will show that the MZW-model can be obtained as a post-Newtonian or large-wrapping decoupling limit of the bosonic membrane.
1.3 A non-relativistic expansion of eleven-dimensional supergravity
The non-relativistic nature of the MZW-model suggests the use of non-relativistic geometry. In particular, we are interested in the application to membranes in eleven-dimensional supergravity. A finite non-relativistic limit of eleven-dimensional supergravity has been obtained in [30, 31] and gives rise to the following expansions of metric $g$ and three-form gauge field $C$ as series of some dimensionless parameter $c\rightarrow\infty$:
$\hat{g}_{\mu\nu}$ $=c^{2}\tau_{\mu}{}^{A}\tau_{\nu}{}^{B}\eta_{AB}+c^{-1}H_{\mu\nu}+\mathcal{O}(c^{-4}),$ (1.5)
$\hat{C}_{\mu_{1}\mu_{2}\mu_{3}}$ $=c^{3}\tau_{\mu_{1}}{}^{A}\tau_{\mu_{2}}{}^{B}\tau_{\mu_{3}}{}^{C}\epsilon_{ABC}+C_{\mu_{1}\mu_{2}\mu_{3}}+\mathcal{O}(c^{-3}).$ (1.6)
$\eta_{AB}$ and $\epsilon_{ABC}$ are the $(1+2)$-dimensional Minkowski metric and totally antisymmetric tensor. The geometry, locally, describes a split into a $(1+2)$-dimensional relativistic and an eight-dimensional non-relativistic space, by three longitudinal (so-called clock) one-formd $\tau^{A}$ and a transverse metric $H$. In inverse of these objects is defined by the completeness and orthogonality conditions
$H^{\mu\rho}H_{\rho\nu}+{\tau^{\mu}}_{A}\tau_{\nu}{}^{A}$ $=\delta_{\nu}^{\mu},$
$H^{\mu\nu}{\tau_{\nu}}^{A}=0$ $=H_{\mu\nu}{\tau^{\nu}}_{A},\quad{\tau^{\mu}}_{A}\tau_{\mu}{}^{B}=\delta_{A}^{B}.$
The relevant examples in this letter are very simplistic, allowing a global split into longitudinal and transverse space with coordinates $x^{\mu}=(x^{M},x^{m})$. Moreover, $\mathrm{d}\tau^{A}=0$, the condition for the so-called Newton-Cartan torsionless membrane geometry, will always be satisfied here.
Similarly to the non-relativistic expansion of the bosonic string [32, 33], a consistent treatment of the membrane in a Newton-Cartan background requires the expansion of the coordinate fields. Here, we assume
$\hat{X}^{\mu}(\sigma)=Y^{\mu}(\sigma)+c^{-3}X^{\mu}(\sigma)+\mathcal{O}(c^{-6}).$ (1.7)
Although this expansion has not been discussed in the literature, [^1] There is only sparse literature on the non-relativistic membrane world-volume. In [34] an $c^{-2}$-expansion for the world-volume fields has been assumed in addition to the $c^{-3}$-expansion of the target space geometry, essentially leading to a resulting expansion at all orders in $c^{-1}$. it coincides with the $c^{-3}$-expansion of the target space and will be crucial in what is to follow.
2 Two decoupling limits for the membrane on T2
In the limit where a bosonic membrane wraps a two-torus $\omega$ times with $\omega\rightarrow\infty$, the effective world-volume coupling scales as $1/\omega$, rendering the theory asymptotically free. This behaviour has been observed in different settings: for example, in de Wit-Hoppe–Nicolai matrix regularisation where the membrane is described by an SU($N$) matrix model with coupling $g\propto T\sim 1/\omega$ [35], or the stretched wrapped membrane setting in [36, 37]. Let us present an incarnation of this setting in the classical membrane world-volume theory.
For this, let us consider the bosonic membrane in a toy model [^2] In the sense, that for now we do not focus on whether this is a solution to eleven-dimensional supergravity. background
$\mathcal{M}_{\tau}\times\mathcal{M}_{H}=(\mathbb{R}_{t}\times T^{2})\times\mathcal{M}_{H}$
supported by a background 4-form flux with on leg in the flat ’longitudinal’ space $\mathcal{M}_{\tau}$ and three in the ’transverse’ space $\mathcal{M}_{H}$.
Here, we present two (ultimately equivalent) views on a decoupling limit that can be interpreted either as a large wrapping limit around the $T^{2}$ or a non-relativistic expansion. In particular, we aim at the following, at leading order in these limits:
  • decoupling limit of the (internal three-dimensional gravity of the) membrane, i.e $\gamma_{MN}\rightarrow\eta_{MN}$
  • membrane dynamics along $\mathcal{M}_{\tau}$ and $\mathcal{M}_{H}$ decouple
  • membrane dynamics along $\mathcal{M}_{H}$rightarrow MZW-model
In the following, we focus on the decoupling of the internal gravity and that, at leading order, the 4-form flux decouples from the $\mathcal{M}_{\tau}$ membrane dynamics. In section 3 it is shown which concrete choice of $\mathcal{M}_{H}$ and $F$ leads to the MZW-model dynamics (1.4) and how this can be embedded into a supergravity solution.
2.1 Post-Newtonian dynamics
Let us propose the following expansion for the background geometry to (2)
$\hat{\mathrm{d}s}^{2}$ $=c^{2}\eta_{MN}\mathrm{d}x^{M}\mathrm{d}x^{N}+c^{-1}H_{mn}\mathrm{d}x^{m}\mathrm{d}x^{n}$ (2.1)
$\hat{F}$ $=F=\frac{1}{3!}F_{klmN}\mathrm{d}x^{k}\wedge\mathrm{d}x^{l}\wedge\mathrm{d}x^{m}\wedge\mathrm{d}x^{N}$ (2.2)
where coordinate $x^{\mu}=(x^{M},x^{m})$ are chosen that make the split into $\mathcal{M}_{\tau}$ and $\mathcal{M}_{H}$ manifest. [^3] Indices of tensor of expansions (without $\hat{}$ are raised and lowered by $\eta$ or $H$. Compared to (1.5), this corresponds to the choice of clock forms $\tau^{M}=\mathrm{d}x^{M}$, which in turn causes the $c^{3}$-contribution to $\hat{F}=\mathrm{d}\hat{C}$ from (1.6) to vanish.
The striven for decoupling of three-dimensional gravity on the membrane world-volume is achieved at non-trivial leading orders of the following expansion for the coordinate fields $\hat{X}^{\mu}(\sigma)$
$\hat{X}^{\mu}(\sigma)$ $=(\hat{X}^{M}(\sigma),\hat{X}^{m}(\sigma))$
$=\left(\sigma^{M}+c^{-3}X^{M}(\sigma),c^{-3}X^{m}(\sigma)\right).$ (2.3)
We assume that $X^{M}(\sigma)$ is periodic in $\sigma^{M}$ for $M=1,2$. The fields $X^{\mu}=(X^{M},X^{m})$ correspond to fluctuations around a static membrane in $\mathbb{R}_{t}\times T^{2}$ that is wrapped around the $T^{2}$. These fluctuations are suppressed by the scale $\epsilon\sim\frac{1}{c^{3}}$. As desired, this expansion decouples the world-volume gravity at leading orders in $\frac{1}{c^{3}}$. The inverse of induced metric $\hat{\gamma}_{MN}=\partial_{M}\hat{X}^{\mu}\partial_{N}\hat{X}^{\nu}\hat{g}_{MN}$ then has the following expansion
$\hat{\gamma}^{MN}=c^{-2}\eta^{MN}-c^{-5}2\partial^{(M}X^{N)}+\mathcal{O}(c^{-8})$ (2.4)
At leading orders indices $M,N,...$ are raised with $\eta$. Moreover, this expansion ensures that the membrane couples (at leading order) only to the leading-order background fields. Moreover, the resulting theory in the decoupling limit is a three-dimensional non-relativistic (non-Lorentz invariant) theory, as is striven for in order to reproduce the MZW-model. The equations motion at leading order are:
00 0 $=\frac{1}{c^{5}}\left(-\partial^{M}\partial_{N}X^{N}+\mathcal{O}\left(c^{-3}\right)\right.$
${}\quad\left.+\frac{1}{c^{6}}\frac{1}{3!}\eta^{MN}F_{Nklm}\partial_{K}X^{k}\partial_{L}X^{l}\partial_{R}X^{m}\epsilon^{KLR}\right)$ (2.5)
00 0 $=\frac{1}{c^{5}}\Big{(}\partial_{M}\partial^{M}X^{m}+...$
${}\quad\left.+\frac{1}{2}H^{mn}F_{klnN}\partial_{K}X^{k}\partial_{L}X^{l}\epsilon^{KLN}+\mathcal{O}\left(c^{-3}\right)\right)$ (2.6)
where ’...’ denotes the Christoffel symbol contribution that will not be relevant in the following. Hence, the four-form flux that mixes longitudinal and transversal degrees of freedom decouples from the longitudinal $X^{M}$ equations of motion.
The leading part of the longitudinal coordinates’ $X^{M}$ dynamics comes from a non-trivial combination the next-to-leading order of the membrane coordinate (2.3) and world-volume metric (2.4) expansions. The resulting powers of $\frac{1}{c}$ in front of the four-form flux are combinations of $\frac{1}{c^{3}}$ factors of the fluctuations $X^{\mu}$ in (2.3) and the effect of raising the index of $F$ with the metric (2.1).
2.2 A large wrapping limit
An alternative point of view can be provided without resorting to membrane Newton-Cartan geometry and non-relativistic interpretation by considering the more general ansatz introducing scales $R$ and $\bar{R}$ for the longitudinal and transversal spaces: [^4] The scaling like this for $F$ has to be introduced in order to be consistent with the supergravity equations. The choice taken here is not general, but the one that will appear below.
$\hat{\mathrm{d}s}^{2}$ $=R^{2}\eta_{MN}\mathrm{d}x^{M}\mathrm{d}x^{N}+\bar{R}^{2}H_{mn}\mathrm{d}x^{m}\mathrm{d}x^{n},$ (2.7)
$\hat{F}$ $=F=\frac{1}{3!}R^{2}\bar{R}F_{klmN}\mathrm{d}x^{k}\wedge\mathrm{d}x^{l}\wedge\mathrm{d}x^{m}\wedge\mathrm{d}x^{N}.$ (2.8)
In contrast to the non-relativistic setting, we look at arbitrary fluctuations around a static wrapped membrane in the large wrapping number limit $\omega\rightarrow\infty$:
$\hat{X}^{\mu}(\sigma)$ $=\left(\hat{X}^{M}(\sigma),\hat{X}^{m}(\sigma)\right)$ (2.9)
$=\left(\omega\sigma^{M}+X^{M}(\sigma),X^{m}(\sigma)\right).$
The parameter $\omega$ describes both a ’world-volume speed of light’ and the wrapping number around the torus. [^5] In principle, these parameters could be different, of course – i.e. $X^{M}=(\omega_{0}\sigma^{0},\omega_{1}\sigma^{1},\omega_{2}\sigma^{2})$. For the limit in question, we want to assume $\omega_{0}\sim\omega_{1}\sim\omega_{2}\sim\omega\rightarrow\infty$. Finite quotients could be compensated for by rescaling of the membrane coordinates $\sigma^{M}$. In the limit $\omega\rightarrow\infty$, the $\mathcal{O}(\omega^{0})$ world-volume fields $x^{\mu}(\sigma)$ are suppressed and the induced metric simplifies significantly. In particular, one can compute the inverse
$\gamma^{MN}=\frac{1}{\omega^{2}R^{2}}\left(\eta^{MN}-\frac{1}{\omega}2\partial^{(M}x^{N)}+\mathcal{O}(\omega^{-2})\right).$ (2.10)
in order to compute the expansion of the equations of motion in leading order
00 0 $=-\frac{1}{\omega^{2}R^{2}}\left(\partial^{K}\partial_{M}X^{M}+\mathcal{O}\left(\frac{1}{\omega}\right)\right)$ (2.11)
${}\quad+\frac{\bar{R}^{2}}{R}\frac{1}{3!}\eta^{KM}F_{Mklm}\partial_{K}X^{k}\partial_{L}X^{l}\partial_{R}X^{m}\epsilon^{KLR},$
00 0 $=\frac{1}{\omega^{2}R^{2}}\left(\partial_{M}\partial^{M}X^{m}+...\right)$ (2.12)
${}\quad+\frac{1}{2}R\eta^{mn}F_{klnN}\partial_{K}X^{k}\partial_{L}X^{l}\epsilon^{KLN}+\mathcal{O}\left(\frac{1}{\omega^{3}R^{2}}\right).$
Requiring that the four-form flux couples to the leading order transverse dynamics and decoupling of the longitudinal and transverse degrees of freedom at leading order leads to the following scaling for the (dimensionless) scales $R$ and $\bar{R}$:
$R^{3}\sim\frac{1}{\omega^{2}}\rightarrow 0,\qquad\mathcal{O}(\bar{R}^{2})<\mathcal{O}\left(\frac{1}{\omega^{2}R}\right).$ (2.13)
The results of the two decoupling limits – the post-Newtonian and the large wrapping one – agree for $\omega\sim c^{3}$, $R\sim c$, and $\bar{R}\sim\frac{1}{\sqrt{c}}$.
3 The MZW-model as decoupling limit of the membrane
3.1 A toy model background
In order to obtain the MZW-model from the transverse equations of motion at leading order, like (2.6), we specify $\mathcal{M}_{H}$ to be a Lie group $G$ [^6] For now, we are not concerned with dimensionality of the background. The background expansion in the conventions for the non-relativstic limit are
$\hat{\mathrm{d}s}^{2}$ $=c^{2}\eta_{MN}\mathrm{d}x^{M}\mathrm{d}x^{N}+c^{-1}\kappa_{ab}e^{a}e^{b}$ (3.1)
$\hat{F}$ $=F=-\frac{1}{3!}f_{abc}e^{a}\wedge e^{b}\wedge e^{c}\wedge v.$
$e^{a}={e_{m}}^{a}\mathrm{d}x^{m}$ are the Maurer-Cartan one-forms on $G$, $f_{abc}$ are the structure constants for the Lie algebra $\mathfrak{g}$ of $G$, and $\kappa$ is the Killing form on $\mathfrak{g}$. $v$ is a one form on $\mathcal{M}_{\tau}$. We choose $v=\mathrm{d}x^{0}$ or $v=\mathrm{d}x^{1/2}$. This choice will be related to the choice of $v$ in the MZW-model (1.4). In the same way as in the previous section, at leading order the membrane equations of motion, now expressed in terms of currents $j^{a}={e_{m}}^{a}\mathrm{d}X^{m}$, are
00 0 $=\partial_{M}\partial_{N}X^{N},$ (3.2)
00 0 $=\partial^{N}j_{N}^{a}-\frac{1}{2}{f^{a}}_{bc}j^{b}_{M}j^{c}_{N}v_{L}\epsilon^{MNL}.´,$ (3.3)
for the expansion (2.3). We recognise the MZW equation of motion (1.4) for the transverse directions. The longitudinal equations of motion $\partial_{M}\partial_{N}X^{N}=0$ decouple, but are solvable in their own right. For the purpose of this paper, we will assume the solution $X^{N}=0$ in the following.
3.2 Embedding into a supergravity solution
[…]
$\mathrm{d}s^{2}$ $=R^{2}\left(\mathrm{d}s^{2}_{\text{AdS}_{3}}+\mathrm{d}s^{2}_{\text{S}^{3}}\right)+\mathrm{d}s^{2}_{T^{4}}+\mathrm{d}z^{2}$
$F$ $=2R^{2}\left(Vol(\text{AdS}_{3})+Vol(\text{S}^{3})\right)\wedge\mathrm{d}z$
where $z$ is the coordinate of the M-theory circle. Moreover, let us single out one of the directions of the $T^{4}$, denoted by $y$, and introduce the scales $R$ and $\bar{R}$
$\mathrm{d}s^{2}$ $=R^{2}\left(\mathrm{d}s^{2}_{\text{AdS}_{3}}+\mathrm{d}s^{2}_{\text{S}^{3}}\right)+\bar{R}^{2}\left(\mathrm{d}y^{2}+\mathrm{d}z^{2}\right)+\mathrm{d}s^{2}_{T^{3}}$
$F$ $=2R^{2}\bar{R}\left(Vol(\text{AdS}_{3})+Vol(\text{S}^{3})\right)\wedge\mathrm{d}z.$ (3.4)
With the following expansion (in a Poincaré patch of AdS3) for example:
$\hat{X}^{\mu}(\sigma)$ $=\left(X_{AdS_{3}}(\sigma);X_{S^{3}}^{m}(\sigma);X_{y}(\sigma),X_{z}(\sigma),X_{T_{3}}(\sigma)\right)$
$X_{AdS_{3}}(\sigma)$ $=(t(\sigma),x(\sigma),r(\sigma))=\left(\omega\frac{\bar{R}}{R}\sigma^{0},0,r_{0}\right)$ (3.5)
$X_{y}(\sigma)$ $=\omega\sigma^{1},\quad X_{z}(\sigma)=\omega\sigma^{2},\quad X_{T^{3}}(\sigma)=const.$
Hence, the focus lies on the transversal fluctuations $X^{m}_{S^{3}}$ around the $S^{3}$, neglecting the other (finally decoupled) fluctuations – both along the remaining transversal directions but also the longitudinal ones. The only significant difference from the toy model case is that the time-like direction lies in AdS3 which scales like the transverse direction. In the above expansion, this is cured by the parameterisation of $t(\sigma)$ including the ration $\frac{\bar{R}}{R}$. Hence, for a large wrapping limit $\omega\rightarrow\infty$, and suitable limits for $R$ and $\bar{R}$, as discussed above (2.13), one reproduces the MZW-model equations of motion for the fluctations $X_{S^{3}}^{m}$.
4 Outlook
In this letter, we proposed a construction that shows that a certain regime of the (bosonic) membrane can be described by a classically integrable theory, the $(1+2)$-dimensional MZW-model, which has been studied in great detail in the literature. In this way, this proposal also offers a physical realisation of this model. Some crucial questions for future research will be:
  • What is the physical interpretation of the conserved charges [22] and integrable hierarchy [24, 38] of the MZW-model for the membrane? Do these also have meaning outside of this regime?
  • What is the physical interpretation of the solitons [39, 40, 25] of the MZW-model on the membrane?
[…]
Moreover, for the integrable two-dimensional $\sigma$-models, generalised geometry resp. O$(d,d)$-duality, was used in order to construct and understand new integrable models [41, 42, 43]. For the membrane a similar formulation, geometrising an underlying $E_{d(d)}/K_{d}$-structure, has been obtained and discussed in the literature [44, 45, 46, 47, 48, 49]. It might be fruitful to systematically study non-relativistic limits and post-Newtonian expansions of this setting in order to obtain new regimes of membrane theory that are described by integrable $(1+2)$-dimensional field theories. In general, the advent of novel higher-dimensional integrable theories via decoupling limits could be an interesting consequence of the present work. As mentioned above, another potential path to new higher-dimensional integrable $\sigma$-models could be possible via the recently suggested higher Chern-Simons theory [29].
Acknowledgements
The author thanks the Institute for Theoretical Physics of ETH Zurich for their warm hospitality when finalising this work and the COST action CA22113 for support under the Short-Term Scientific Mission grant E-COST-GRANT-CA22113-2a51a346.