### v2

# The Manakov-Zakharov-Ward model as an integrable decoupling limit of the membrane

Abstract
[…]
1 Introduction
[…]
In addition to these previously mentioned understood corners of membrane theory, this letter suggests a novel tractable regime – via a decoupling limit of the membrane in a certain background geometry – described by a $(1+2)$-dimensional classically integrable theory. Integrable field theories in $(1+1)$ dimensions evade typical no-go constraints by admitting infinitely many local conserved charges that render scattering purely elastic and factorised. However, in $(1+2)$ or higher dimensions, the same infinite symmetries force the S-matrix to be trivial (that is, free), as any non-trivial scattering would violate the independence from impact parameters. Thus, beyond one spatial dimension, integrable models must break Lorentz invariance. One of the few existing examples of such non-Lorentz invariant integrable theories is the Manakov-Zakharov-Ward model [7, 8].
A non-Lorentzian decoupling limit, as proposed here, fits into a recently observed pattern that certain decoupling limits that isolate BPS pp‑brane dynamics and associated matrix models [9, 10, 11], naturally giving rise to non‑Lorentzian target‑space geometries known as $p$-brane Newton-Cartan geometry (see [12] for a recent review). More recently, this has also been discussed extended to eleven-dimensional supergravity [13, 14, 15, 16].
1.1 The bosonic membrane
The action of the bosonic membrane (or equivalently the bosonic part of the supermembrane [2]) is
$S$ $\sim\int\mathrm{d}^{3}\sigma\left(\frac{1}{2}\sqrt{-\gamma}\gamma^{MN}\partial_{M}\hat{X}^{\mu}\partial_{N}\hat{X}^{\nu}g_{\mu\nu}\right$. (1.1)
${}\qquad+\left.\frac{1}{3!}\epsilon^{KLM}\partial_{K}\hat{X}^{\mu}\partial_{L}\hat{X}^{\nu}\partial_{M}\hat{X}^{\rho}C_{\mu\nu\rho}-\frac{1}{2}\sqrt{-\gamma}\right)$
in a bosonic supergravity background described by a metric $g$ and three-form gauge field $C$. $\sigma^{M}$ are the coordinates on the membrane world-volume, $\hat{X}^{\mu}(\sigma)$ the embedding coordinate fields, and $\gamma_{MN}$ is the (dynamical) world-volume metric. The cosmological constant term, as usual, ensures that the world-volume metric becomes the induced metric (as algebraic equation of motion for the world-volume metric $\gamma$):
$\gamma_{MN}=\partial_{M}\hat{X}^{\mu}\partial_{N}\hat{X}^{\nu}g_{\mu\nu}.$ (1.2)
In contrast to the string, where reparameterisations and Weyl invariance can be used to gauge-fix the world-volume metric to be flat, this is not the case for the membrane. The equations of motion for the world-volume coordinate fields $\hat{X}^{\mu}(\sigma)$ are
$0$ $=\partial_{M}(\sqrt{-\gamma}\gamma^{MN}\partial_{N}\hat{X}^{\mu})+\sqrt{-\gamma}\gamma^{MN}\Gamma^{\mu}{}_{\kappa\lambda}\partial_{M}\hat{X}^{\kappa}\partial_{N}\hat{X}^{\lambda},$
${}\quad+\epsilon^{NKL}\frac{1}{3!}{F^{\mu}}_{\nu\kappa\lambda}\partial_{N}\hat{X}^{\nu}\partial_{K}\hat{X}^{\kappa}\partial_{L}\hat{X}^{\lambda},$ (1.3)
where $\Gamma$ denote Christoffel symbols and $F=\mathrm{d}C$. Due to the non-trivial nature of the world-volume metric $\gamma$, these equations are quite intractable, in general.
Famously [17, 18], even this supposedly most simple case leads to complications, such as a continuous spectrum and spike instabilities. Quantum non-integrability of the membrane in the sense of scattering of fluctuations around a classical solution, and its relation to string theory, has recently been shown in [19]. In this letter, we will see that a large wrapping limit together with a scaling limit of the background can simplify these equations. In the special case considered here, the result will even be a classically integrable system.
1.2 The Manakov-Zakharov-Ward model
Classically integrable field theories in dimensions higher than $(1+1)$ are rare, especially those of the $\sigma$-model type, which are abundant in $(1+1)$ dimensions (see [20, 21] for recent reviews). A notable exception is the Manakov-Zakharov-Ward (MZW) model [7, 8], an integrable field theory in $(1+2)$ dimensions. It arises via dimensional reduction from the anti-self-dual Yang-Mills equations in $(2+2)$-dimensional spacetime. This reduction involves assuming that the gauge potentials are independent of one spatial coordinate, resulting in a non-Lorentz invariant system in $(1+2)$ dimensions. [^1] As a recent approach to obtain new integrable models in higher dimensions, let us mention [22].
The model is of $\sigma$-model type and can be formulated in terms of maps $g:\Sigma\rightarrow G$ from a $(1+2)$-dimensional flat spacetime $\Sigma$ (referred to as the world-volume) into a Lie group $G$. The dynamics is conveniently expressed using the left-invariant Maurer–Cartan current $j=g^{-1}\mathrm{d}g\in\mathfrak{g}$, valued in the Lie algebra $\mathfrak{g}$ of $G$. The flatness condition $\mathrm{d}j+\frac{1}{2}[j,j]=0$ is automatically satisfied. The equations of motion take the form
$\mathrm{d}\star j+\mathrm{d}j\wedge v=\mathrm{d}\star j-\frac{1}{2}[j,j]\wedge v=0,$ (1.4)
where $v$ is a constant world-volume vector that explicitly breaks Lorentz invariance in $(1+2)$ dimensions.
Originally introduced by Manakov and Zakharov [7] for the case of $G=\mathrm{SU}(2)$, the model was shown to be solvable by the inverse scattering transform. It supports soliton solutions localised in all spatial directions and does not show classical scattering, an identifying feature of integrability [8]. Ward [8] further demonstrated that for the model to have a well-defined energy functional, the vector $v$ must be space-like, although its integrability remains unaffected by the nature of $v$. The MZW-model possesses an infinite hierarchy of conserved charges arising from its Lax pair and zero-curvature formulation [8]. Although only the energy-momentum tensor yields local, finite Noether charges, Ioannidou and Ward constructed an infinite sequence of non-local but finite conserved charges using integrals of the currents and its powers [23]. These charges reflect the underlying loop-algebra and twistor-theoretic structure of the model [24], and parallel the hierarchy of integrable flows found in systems such as the matrix KP-hierarchy [25]. This integrable structure also governs soliton interactions and scattering phenomena [26].
Recently, the MZW-model has reemerged in the context of higher-dimensional Chern–Simons theory [27, 28], a unifying framework for integrable $\sigma$-models associated with Lie group symmetries. In this setting, the MZW-model appears as a special case in certain higher-dimensional generalisations [29, 30]. Partly because of its breaking of Lorentz invariance, it was not obvious in which physical settings the MZW-model appears. Moreover, it seemed unlikely that the MZW-model will play a role in membrane dynamics since it assumes a flat metric on $\Sigma$ in contrast to the dynamical induced metric of the bosonic membrane. However, this article will show that the MZW-model can be obtained as a post-Newtonian or large-wrapping decoupling limit of the bosonic membrane.
2 A decoupling limit for the membrane on T2
In the limit where a bosonic membrane wraps a two-torus $\omega$ times with $\omega\rightarrow\infty$, the effective world-volume coupling scales as $1/\omega$, rendering the theory asymptotically free. This behaviour has been observed in different settings: for example, in de Wit-Hoppe–Nicolai matrix regularisation where the membrane is described by an SU($N$) matrix model with coupling $g\propto T\sim 1/\omega$ [31], or the stretched wrapped membrane setting in [32, 33]. Let us present an incarnation of this setting in the classical membrane world-volume theory.
For this, let us consider the bosonic membrane in a toy model [^2] In the sense, that for now we do not focus on whether this is a solution to eleven-dimensional supergravity. background
$\mathcal{M}_{\tau}\times\mathcal{M}_{H}=(\mathbb{R}_{t}\times T^{2})\times\mathcal{M}_{H}$
supported by a background 4-form flux with on leg in the flat ’longitudinal’ space $\mathcal{M}_{\tau}$ and three in the ’transverse’ space $\mathcal{M}_{H}$.
Here, we present two (ultimately equivalent) views on a decoupling limit that can be interpreted either as a large wrapping limit around the $T^{2}$ or a non-relativistic expansion (presented in appendix A). In particular, we aim at the following, at leading order in these limits:
  • decoupling limit of the (internal three-dimensional gravity of the) membrane, i.e $\gamma_{MN}\rightarrow\eta_{MN}$
  • membrane dynamics along $\mathcal{M}_{\tau}$ and $\mathcal{M}_{H}$ decouple
  • membrane dynamics along $\mathcal{M}_{H}$ $\rightarrow$ MZW-model
This will involve scaling limits – both on the target space and one on the membrane world-volume.
In the following, we focus on the decoupling of the internal gravity and that, at leading order, the 4-form flux decouples from the $\mathcal{M}_{\tau}$ membrane dynamics. In section 3 it is shown which concrete choice of $\mathcal{M}_{H}$ and $F$ leads to the MZW-model dynamics (1.4) and how this can be embedded into a supergravity solution.
For this, let consider a general ansatz introducing scales $R$ and $\bar{R}$ for the longitudinal ($\mathcal{M}_{\tau}$, coordinates [^3] In a slight abuse of notation, we take the same indices $K,L,M,$... for the transversal coordinates $\hat{X}^{M}$ and the world-volume coordinates $\sigma^{M}$. $\hat{X}^{M}$) and transversal spaces ($\mathcal{M}_{H}$, coordinates $\hat{X}^{m}$)
$\mathrm{d}s^{2}$ $=R^{2}\eta_{MN}\mathrm{d}\hat{X}^{M}\mathrm{d}\hat{X}^{N}+\bar{R}^{2}H_{mn}\mathrm{d}\hat{X}^{m}\mathrm{d}\hat{X}^{n},$ (2.1)
$F$ $=\frac{1}{3!}R\bar{R}^{2}F_{klmN}\mathrm{d}\hat{X}^{k}\wedge\mathrm{d}\hat{X}^{l}\wedge\mathrm{d}\hat{X}^{m}\wedge\mathrm{d}\hat{X}^{N}.$ (2.2)
The scaling of $F$ has to be introduced in order to be consistent with the supergravity equations – compare (3.3).
Moreover, on the world-volume we consider arbitrary fluctuations around a static wrapped membrane in a large wrapping number limit $\omega\rightarrow\infty$:
$\hat{X}^{\mu}(\sigma)$ $=\left(\hat{X}^{M}(\sigma),\hat{X}^{m}(\sigma)\right)$ (2.3)
$=\left(\omega\sigma^{M}+X^{M}(\sigma),X^{m}(\sigma)\right).$
The parameter $\omega$ describes both a ’world-volume speed of light’ and the wrapping number around the torus. [^4] In principle, these parameters could be different, of course – i.e. $X^{M}=(\omega_{0}\sigma^{0},\omega_{1}\sigma^{1},\omega_{2}\sigma^{2})$. For the limit in question, we want to assume $\omega_{0}\sim\omega_{1}\sim\omega_{2}\sim\omega\rightarrow\infty$. Finite quotients could be compensated for by rescaling of the membrane coordinates $\sigma^{M}$. In the limit $\omega\rightarrow\infty$, the $\mathcal{O}(\omega^{0})$ world-volume fields $X^{\mu}(\sigma)$ are suppressed and the induced metric simplifies significantly. In particular, one can compute
$\gamma^{MN}$ $=\frac{1}{\omega^{2}R^{2}}\left(\eta^{MN}-\frac{1}{\omega}2\partial^{(M}X^{N)}+\mathcal{O}(\omega^{-2})\right),$ (2.4)
$\sqrt{-\gamma}$ $=\omega^{3}R^{3}\left(1+\frac{1}{\omega}\partial_{M}X^{M}+\mathcal{O}(\omega^{-2})\right),$
in order to compute the expansion of the equations of motion (1.3) for $X^{M}$ and $X^{M}$ at leading order
$0$ $=\omega R\left(0+\mathcal{O}\left(\omega^{-1}\right)\right)$ (2.5)
${}\quad+\frac{\bar{R}^{2}}{R}\frac{1}{3!}\eta^{KM}F_{Mklm}\partial_{N}X^{k}\partial_{L}X^{l}\partial_{R}X^{m}\epsilon^{NLR},$
[…]
${}\quad\left.+\frac{1}{2}H^{mn}F_{klnN}\partial_{K}X^{k}\partial_{L}X^{l}\epsilon^{KLN}+\mathcal{O}\left(\omega^{-1}\right)\right).$
Besides $\omega\rightarrow\infty$, the following scaling for the (dimensionless) scales $R$ and $\bar{R}$ is required, simultaneously ensuring the form (2.4) and guaranteeing that the transversal fluctations decouple from from the four-form flux at leading order:
$\frac{\bar{R}^{2}}{R^{2}\omega}\rightarrow 0.$ (2.7)
Remarkably, at leading order in the large wrapping limit the contributions from the metric fluctations (2.4) cancel completely in the equations motion (2.5).
There are several ways to interpret this limit. Let us outline two interpretations with very different consequences:
  • One can solve (2.7) with the non-Lorentzian parameterisation introduced in appendix A. For this let us introduce the dimensionless scale $c$ with
    $R^{2}=c^{2},\qquad\bar{R}^{2}=\frac{1}{c}.$ (2.8)
    In order to satisfy (2.7), the scaling of $\omega\sim c^{\alpha}\rightarrow\infty$ has to be chosen accordingly in terms of $c$. There are two choices
    • a non-relativistic limit $c\rightarrow\infty$: $\alpha>0$.
    • a hyper-relativistic (Carollian) limit $c\rightarrow 0$: $\alpha<-3$.
    Interestingly, in both cases this corresponds to large energy limits $E=\omega R\rightarrow\infty$. Although at leading order one reproduces the MZW-model, these cases will lead to additional equations of motion at the finite but subleading order. [^5] Additional contributions would arise at order $\mathcal{O}(R)<\mathcal{O}(\omega R)$ or $\mathcal{O}(\bar{R}^{2}/R)\mathcal{O}(\omega R)$.
  • One can require that all the subleading contributions are supressed, and only the leading order contributions (proportional to $\omega R$) could be potentially finite. This leads to the following additional set of constraints.
    $R\rightarrow 0,\quad\frac{\bar{R}^{2}}{R}\rightarrow 0$ (2.9)
    This is solved by a parameterisation $R\sim\bar{R}\sim\frac{1}{\omega^{\alpha}}$, with $\alpha>0$. Then this limit can be interpreted as either a large or small energy $\omega R$ limit for $\alpha>0$ or $\alpha<0$ respectively. In any case, this is a small volume regime in target space – both scales $R,\bar{R}\rightarrow 0$.
3 The MZW-model as decoupling limit of the membrane
3.1 A toy model background
In order to obtain the MZW-model from the transverse equations of motion at leading order, like (2.6), we specify $\mathcal{M}_{H}$ to be a Lie group $G$.
$\mathrm{d}s^{2}$ $=R^{2}\eta_{MN}\mathrm{d}\hat{X}^{M}\mathrm{d}\hat{X}^{N}+\bar{R}^{2}\kappa_{ab}e^{a}e^{b}$ (3.1)
$F$ $=-R\bar{R}^{2}\frac{1}{3!}f_{abc}e^{a}\wedge e^{b}\wedge e^{c}\wedge v.$
$e^{a}={e_{m}}^{a}\mathrm{d}\hat{X}^{m}$ are the Maurer-Cartan one-forms on $G$, $f_{abc}$ are the structure constants for the Lie algebra $\mathfrak{g}$ of $G$, and $\kappa$ is the Killing form on $\mathfrak{g}$. $v$ is a one form on $\mathcal{M}_{\tau}$. We choose $v=\mathrm{d}\hat{X}^{0}$ or $v=\mathrm{d}\hat{X}^{i}$, for $i=1,2$. This choice will be related to the choice of $v$ in the MZW-model (1.4). In the same way as in the previous section, at leading order the membrane equations of motion, now expressed in terms of currents $j^{a}={e_{m}}^{a}\mathrm{d}X^{m}$, are
$0$ $=\partial^{N}j_{N}^{a}-\frac{1}{2}{f^{a}}_{bc}j^{b}_{M}j^{c}_{N}v_{L}\epsilon^{MNL},$ (3.2)
for the expansion (2.3). We recognise the MZW-model’s equation of motion (1.4) for the transverse directions. The longitudinal equations of motion completely decouples at leading order as in (2.5).
3.2 Embedding into a supergravity solution
[…]
$\mathrm{d}s^{2}$ $=\bar{R}^{2}\left(\mathrm{d}s^{2}_{\text{AdS}_{3}}+\mathrm{d}s^{2}_{\text{S}^{3}}\right)+\mathrm{d}s^{2}_{T^{4}}+\mathrm{d}z^{2}$
$F$ $=2\bar{R}^{2}\left(Vol(\text{AdS}_{3})+Vol(\text{S}^{3})\right)\wedge\mathrm{d}z$
where $z$ is the coordinate of the M-theory circle. Moreover, let us single out one of the directions of the $T^{4}$, denoted by $y$, and introduce the scales $R$ and $\bar{R}$
$\mathrm{d}s^{2}$ $=\bar{R}^{2}\left(\mathrm{d}s^{2}_{\text{AdS}_{3}}+\mathrm{d}s^{2}_{\text{S}^{3}}\right)+R^{2}\left(\mathrm{d}y^{2}+\mathrm{d}z^{2}\right)+\mathrm{d}s^{2}_{T^{3}}$
$F$ $=2R\bar{R}^{2}\left(Vol(\text{AdS}_{3})+Vol(\text{S}^{3})\right)\wedge\mathrm{d}z.$ (3.3)
With the following expansion (in a Poincaré patch of AdS3) for example:
$\hat{X}^{\mu}(\sigma)$ $=\left(X_{AdS_{3}}(\sigma);X_{S^{3}}^{m}(\sigma);X_{y}(\sigma),X_{z}(\sigma),X_{T_{3}}(\sigma)\right)$
$X_{AdS_{3}}(\sigma)$ $=(t(\sigma),x(\sigma),r(\sigma))=\left(\omega\frac{R}{\bar{R}}\sigma^{0},0,r_{0}\right)$ (3.4)
$X_{y}(\sigma)$ $=\omega\sigma^{1},\quad X_{z}(\sigma)=\omega\sigma^{2},\quad X_{T^{3}}(\sigma)=const.$
Hence, the focus lies on the transversal fluctuations $X^{m}_{S^{3}}$ around the $S^{3}$, neglecting the other (finally decoupled) fluctuations – both along the remaining transversal directions but also the longitudinal ones. The only significant difference from the toy model case is that the time-like direction lies in AdS3 which scales like the transverse direction. In the above expansion, this is cured by the parameterisation of $t(\sigma)$ including the ratio $\frac{R}{\bar{R}}$. Hence, for a large wrapping limit $\omega\rightarrow\infty$, and suitable limits for $R$ and $\bar{R}$, as discussed above (2.7), one reproduces the MZW-model equations of motion for the fluctations $X_{S^{3}}^{m}$.
4 Outlook
In this letter, we proposed a construction that shows that a certain regime of the (bosonic) membrane can be described by a classically integrable theory, the $(1+2)$-dimensional MZW-model, which has been studied in great detail in the literature. In this way, this proposal also offers a physical realisation of this model. Some crucial questions for future research will be:
  • What is the physical interpretation of the conserved charges [23] and integrable hierarchy [25, 34] of the MZW-model for the membrane? Do these also have meaning outside of this regime?
  • What is the physical interpretation of the solitons [35, 36, 26] of the MZW-model on the membrane?
[…]
Moreover, for the integrable two-dimensional $\sigma$-models, generalised geometry resp. O(d,d)(d,d)-duality, was used in order to construct and understand new integrable models [37, 38, 39]. For the membrane a similar formulation, geometrising an underlying $E_{d(d)}$-structure, has been obtained and discussed in the literature [40, 41, 42, 43, 44, 45]. It might be fruitful to systematically study non-Lorentzian expansions of this setting in order to obtain new regimes of membrane theory that are described by integrable $(1+2)$-dimensional field theories. In general, the advent of novel higher-dimensional integrable theories via decoupling limits could be an interesting consequence of the present work. As mentioned above, another potential path to new higher-dimensional integrable $\sigma$-models could be possible via the recently suggested higher Chern-Simons theory [30].
Acknowledgements
The research of D.O. is part of project No. 2024/55/D/ST2/01205 funded by the National Science Centre (NCN) of Poland. The author thanks the Institute for Theoretical Physics of ETH Zurich for their warm hospitality when finalising this work, the participants of the ’Workshop on Higher-d Integrability’ in Favignana for fruitful discussions about the draft, and the COST action CA22113 for support under the Short-Term Scientific Mission grant E-COST-GRANT-CA22113-2a51a346. Special credits belong to the anonymous referee of Physical Review Letters pointing out a crucial mistake in the original draft.
Appendix A The non-Lorentzian expansion of supergravity
The non-Lorentzian nature of the MZW-model suggests the use of non-Lorentzian geometry. In particular, we are interested in the application to membranes in eleven-dimensional supergravity. A finite non-relativistic limit of eleven-dimensional supergravity has been obtained in [46, 47] and gives rise to the following expansions of metric $g$ and three-form gauge field $C$ as series of some dimensionless parameter:
$g_{\mu\nu}$ $=c^{2}\tau_{\mu}{}^{A}\tau_{\nu}{}^{B}\eta_{AB}+c^{-1}H_{\mu\nu},$ (A.1)
$C_{\mu_{1}\mu_{2}\mu_{3}}$ $=c^{3}\tau_{\mu_{1}}{}^{A}\tau_{\mu_{2}}{}^{B}\tau_{\mu_{3}}{}^{C}\epsilon_{ABC}+C_{\mu_{1}\mu_{2}\mu_{3}}.$ (A.2)
Typically, the non-relativistic limit $c\rightarrow\infty$ was considered in the 11d supergravity case, but in other cases also the hyper-relativistic or Carollian limit ($c\rightarrow 0$) was considered from similar expansions in general relavitiy [48]. Both cases are of interest in main text.
$\eta_{AB}$ and $\epsilon_{ABC}$ are the $(1+2)$-dimensional Minkowski metric and totally antisymmetric tensor. The geometry, locally, describes a split into a $(1+2)$-dimensional relativistic and an eight-dimensional non-relativistic space, by three longitudinal (so-called clock) one-forms $\tau^{A}$ and a transverse metric $H$. In inverse of these objects is defined by the completeness and orthogonality conditions
$H^{\mu\rho}H_{\rho\nu}+{\tau^{\mu}}_{A}\tau_{\nu}{}^{A}$ $=\delta_{\nu}^{\mu},$
$H^{\mu\nu}{\tau_{\nu}}^{A}=0$ $=H_{\mu\nu}{\tau^{\nu}}_{A},\quad{\tau^{\mu}}_{A}\tau_{\mu}{}^{B}=\delta_{A}^{B}.$
The relevant examples in this letter are very simplistic, allowing a global split into longitudinal and transverse space with coordinates $\hat{X}^{\mu}=(\hat{X}^{M},\hat{X}^{m})$. Moreover, $\mathrm{d}\tau^{A}=0$, the condition for the so-called Newton-Cartan torsionless membrane geometry, will always be satisfied here.